# A Comprehensive Evaluation of Large Language Models on Mental Illnesses

### v1

@@ Abstract
Large Language Model (LLM) s have shown promise in various domains, including healthcare. In this study, we conducted a comprehensive evaluation of LLM s in the context of mental health tasks using social media data. We explored the zero-shot (ZS) and few-shot (FS) capabilities of various LLM s, including GPT-4, Llama 3, Claude, Gemma, Gemini, Phi-3, and others, on tasks such as binary disorder detection, disorder severity evaluation, and psychiatric knowledge assessment. Our evaluation involved 33 models ranging from 2 billion to 405+ billion parameters, and we utilized 9 main prompt templates across the tasks. Key findings revealed that models like GPT-4 and Llama 3 exhibited superior performance in binary disorder detection, with accuracies reaching up to 85% on certain datasets. Moreover, prompt engineering played a crucial role in enhancing model performance. For example, when transitioning from our first binary prompt to a more structured prompt, the accuracy of certain models increased significantly. Notably, the Mixtral 8x22b model showed an improvement of over 20%, while Gemma 7b experienced a similar boost in performance. In the task of disorder severity evaluation, we observed that FS learning significantly improved the model’s accuracy, highlighting the importance of contextual examples in complex assessments. Notably, the Phi-3-mini model exhibited a substantial increase in performance, with balanced accuracy (BA) improving by over 6.80% and mean average error (MAE) dropping by nearly 1.3 when moving from ZS to FS learning. Additionally, in the psychiatric knowledge assessment task, recent models generally outperformed older, larger counterparts, with the Llama 3.1 405b achieving an accuracy of 91.2%. Despite promising results, our analysis identified several challenges, including variability in performance across datasets and the need for careful prompt engineering. There is also the high cost associated with using large models and the limitations imposed by the quality of social media data. Furthermore, the ethical guards imposed by many LLM providers hamper the ability to accurately evaluate their performance, due to tendency to not respond to potentially sensitive queries.
@@ 1 Introduction
Artificial Intelligence (AI) is rapidly transforming the landscape of mental healthcare, offering innovative solutions to address the growing global burden of mental illness. According to the World Health Organization, mental disorders accounted for 5.1% of the global disease burden in 2019, with depressive disorders affecting 280 million people worldwide, and an estimated 703,000 people dead by suicide [1]. The economic consequences of mental health conditions are also significant. In the United States, the cumulative cost of mental health inequities is projected to reach a staggering US\$14 trillion between 2024 and 2040. These costs include direct medical expenses, emergency department utilization, productivity losses, and premature deaths [2].
@@
The integration of AI in psychiatry has enabled a wide array of applications, spanning from the crucial tasks of early detection and diagnosis to predicting treatment outcomes and providing therapeutic interventions. AI models have been employed to analyze diverse data modalities, including neuroimaging data such as functional magnetic resonance imaging (fMRI) and electroencephalogram (EEG), physiological signals like electrocardiogram (ECG), genetic information, and demographic data [3].
@@
The advent of Natural Language Processing (NLP) in mental health research opened up new possibilities. Early NLP applications focused on analyzing text from clinical interviews and self-report questionnaires. These applications used techniques like the Linguistic Inquiry and Word Count to quantify linguistic patterns associated with various mental health conditions [4]. The emergence of transformer-based models, such as Bidirectional Encoder Representations from Transformers (BERT) [5], marked a significant turning point in NLP for mental health. These models, trained on massive amounts of text data, can automatically learn complex linguistic representations, eliminating the need for manual feature engineering. This has led to substantial improvements in tasks like sentiment analysis, suicide risk assessment, and the detection of mental health conditions from social media posts [6].
@@
The advancement to LLM s marks a significant leap forward in the field of NLP. LLM s, such as OpenAI’s GPT-4 [7], Google’s Gemini [8, 9], Meta’s Llama [10], and many others have been trained on vast corpora of text data from diverse sources across the internet, encompassing billions of words and a wide range of human knowledge. This extensive training enables LLM s to understand and generate human-like text with a remarkable degree of fluency and coherence.
@@
Given their extensive training, LLM s have immense potential in the realm of mental health and psychiatry. They can be employed for initial consultations, providing patients with preliminary information about mental health disorders and guiding them toward appropriate care pathways. LLM s can quickly analyze transcripts from patient sessions, offering initial insights and supporting medical practitioners in making more informed decisions. Their ability to process large volumes of patient history and previous medication data can streamline administrative tasks, ensuring that practitioners have quick access to comprehensive patient profiles. Additionally, LLM s can serve as virtual counselors, offering support and guidance to individuals experiencing poor mental states, thus extending the reach of mental health services to those who may not have immediate access to professional help. These models can also assist in monitoring social media for signs of mental distress, enabling timely interventions and potentially saving lives. As these technologies continue to evolve, their integration into mental healthcare promises to enhance the efficiency, accessibility, and quality of care provided to patients [11].
However, most if not all of these capabilities have yet to be properly verified for use in any real medical context. While the potential is significant, rigorous validation and clinical trials are necessary to ensure their effectiveness and safety in practical applications. Current work in the literature revolves around a few key topics, including evaluation (Sec. 1.1), fine-tuning (Sec. 1.2), data augmentation and chatbots (Sec. 1.3), benchmarks (Sec. 1.4), and literature reviews (Sec. 1.5).
@@ 1.1 Evaluating LLM s on Mental Health Tasks
One of the earliest works in this domain is the study by [12], which presents a simple evaluation of GPT-3.5-Turbo’s ZS binary classification performance on stress, depression, and suicidality detection tasks using social media posts. The resulting F1 scores (73% for stress, 86% for depression, and 37% for suicide), were particularly bad in the suicidality severity prediction task. The authors of [13] extend this evaluation by assessing GPT-3.5-Turbo’s text classification abilities on a broader range of affective computing problems, including big-five personality prediction and sentiment analysis, in addition to suicide tendency detection. Their findings reveal that while GPT-3.5-Turbo provided decent results, comparable to Word2Vec and bag-of-words baselines, it did not outperform the fine-tuned RoBERTa model on the given tasks. However, the authors did not prompt GPT-3.5-Turbo through an Application Programming Interface (API), they instead did it manually using the web interface, only resetting the context every 25 prompts, which might have affected their results.
@@ 1.2 Fine-tuning LLM s for mental health tasks
Building upon the evaluation studies, [14] and [15] explore fine-tuning LLM s for improved performance in mental health tasks. [14] introduce the MentaLlama model, which is fine-tuned on the IMHI dataset, a large dataset of social media posts labeled with mental health conditions and explanations. The authors created the IMHI dataset by taking multiple existing mental health datasets and generating additional explanatory labels using GPT-3.5-Turbo. The MentaLlama-chat-13B model surpasses or approaches state-of-the-art discriminative methods in prediction correctness on 7 out of 10 test sets and generates explanations on par with GPT-3.5-Turbo thus providing interpretability.
The authors of [15] conducted a comprehensive evaluation of multiple LLM s, including Alpaca, Alpaca-LoRA, FLAN-T5, Llama 2, GPT-3.5-Turbo, and GPT-4, on various mental health prediction tasks using online text data. Their work explored ZS prompting, FS prompting, and instruction fine-tuning techniques. The authors fine-tuned Alpaca and FLAN-T5 on six mental health prediction tasks across four datasets from Reddit. Results show that instruction fine-tuning significantly improves the performance of LLM s on these tasks. The fine-tuned models, Mental-Alpaca and Mental-FLAN-T5, outperformed much larger models like GPT-3.5-Turbo and GPT-4 on specific tasks, and performed on par with the state-of-the-art task-specific model Mental-RoBERTa [6]. The study also includes an exploratory case study on these models’ reasoning capability, which further suggests both the promising future and the important limitations of LLM s.
@@ 1.3 LLM s for Data Augmentation and Chatbot Development
[16], [17], and [18] explore the application of LLM s beyond prediction tasks, focusing on data augmentation and chatbot development. [16] develop Chat-Diagnose, an explainable and interactive LLM-augmented system for depression detection in social media. Chat-Diagnose incorporates a tweet selector to filter excessive posts, an image descriptor to convert images into text, and professional diagnostic criteria (Diagnostic and Statistical Manual of Mental Disorders (DSM)-5) to guide the diagnostic process. The system utilizes the chain-of-thought (CoT) technique to provide explanations and diagnostic evidence. In the full-training setting, it leverages answer heuristics from a traditional depression detection model. Chat-Diagnose achieves state-of-the-art (SoTA) performance in various settings, including ZS, FS, and full-training scenarios.
[17] investigates the use of GPT-3.5-Turbo for generating synthetic Reddit posts simulating depression symptoms from the Beck Depression Inventory (BDI)-II questionnaire, aiming to enhance semantic search capabilities. The study finds that using original BDI-II responses as queries is more effective, suggesting that the generated data might be too specific for this task.
[18] focuses on developing chatbots that simulate psychiatrists and patients in clinical diagnosis scenarios, specifically for depressive disorders. The study follows a three-phase design, involving collaboration with psychiatrists, experimental studies, and evaluations with real psychiatrists and patients. The development process emphasizes an iterative refinement of prompt design and evaluation metrics based on feedback from both psychiatrists and patients. A key finding of the study is the comparison between real and simulated psychiatrists, revealing differences in questioning strategies and empathy behaviors.
@@ 1.4 Benchmarks for Evaluating LLM s in Psychiatry
In contrast to the aforementioned studies, the authors of [19] took a different approach. Instead of evaluating existing LLM s on established datasets, they sought to create a comprehensive benchmark specifically designed to assess the capabilities of LLM s in the mental health domain. PsyEval (the benchmark) includes tasks such as mental health question-answering, where LLM s are evaluated on their ability to provide accurate and informative responses to queries related to mental health conditions, symptoms, and treatments. The benchmark also features tasks focused on diagnosis prediction, both using data from social media or simulated dialouges between patients and clinicians. Additionally, PsyEval incorporates tasks that assess LLM s’ ability to provide empathetic and safe psychological counseling in simulated conversations.
@@ 1.5 Literature Reviews
In addition to the contribution papers mentioned above, several surveys have been conducted to summarize and analyze the landscape of LLM s in mental health care. The authors of [11] provide a comprehensive overview of the opportunities and risks associated with LLM s in psychiatry, discussing their potential to enhance mental health care through improved diagnostic accuracy, personalized care, and streamlined administrative processes. The authors highlight the ability of LLM s to efficiently analyze patient data, summarize therapy sessions, and aid in complex diagnostic problem-solving. They also discuss the potential of LLM s to automate certain managerial decisions within hospital systems, such as personnel schedules and equipment needs. However, the authors also acknowledge the risks associated with LLM s, such as labor substitution, the potential for reduced human-to-human socialization, and the amplification of existing biases. They conclude by advocating for the development of pragmatic frameworks, including red-teaming and multi-stakeholder-oriented safety and ethical guidelines, to ensure the safe and responsible deployment of LLM s in psychiatry.
@@
The authors of [20] conducted a systematic review of the applications of LLM s, such as GPT-3.5-Turbo, in the field of psychiatry. The review identified 16 studies that directly examined the use of LLM s in psychiatry, with applications ranging from clinical reasoning and social media analysis to education. The authors found that LLM s like GPT-3.5-Turbo and GPT-4 showed promise in assisting with tasks such as diagnosing mental health issues, managing depression, evaluating suicide risk, and supporting education in the field. However, the review also highlighted limitations, such as difficulties with complex cases and potential underestimation of suicide risks.
@@
The article by [21] presents a scoping review of LLM s in mental health care, focusing on their applications and outcomes. The review identified 34 publications that met their inclusion criteria, including studies compiling new datasets, studies focused on model development, studies focused on employing and evaluating LLM s, and studies examining the potential opportunities, risks and ethical considerations associated with utilizing LLM s in a psychiatric context. The authors provide a detailed overview of the models and datasets used in these studies. They found that most studies utilized GPT-3.5 or GPT-4, with some fine-tuning on mental health-specific datasets like MentaLlama and Mental-LLM. The datasets used in these studies were primarily sourced from social media platforms and online mental health forums, with a few utilizing clinical data or simulated conversations. The authors also discuss the key challenges associated with LLM s, such as data availability and reliability, effective handling of mental states, and appropriate evaluation methods.
1.6 Research Gaps
While the potential of LLM s in mental health care is evident, several research gaps remain to be addressed. Firstly, most studies have employed relatively outdated models like GPT-3.5-Turbo, GPT-4, and Llama 2, with no investigations into newer models such as GPT-4-0, GPT-4-Turbo, Llama 3, Phi-3, Gemma 1 & 2, Claude models, and others. This limits our understanding of how the latest advancements in LLM technology can be leveraged for mental health applications. Secondly, there is a lack of focus on prompt engineering to investigate the effect of different prompts on disorder detection tasks. While a few studies have touched upon this aspect, a more systematic and comprehensive exploration of prompt engineering techniques is needed to optimize LLM performance in this domain. Thirdly, many studies have overlooked smaller yet important factors such as LLM variation across runs and the effect of small prompt modifications across models. Additionally, there is often a lack of full transparency in reporting methodologies, even when code is provided, hindering reproducibility and further research.
The goal of this work is to comprehensively evaluate the capabilities of LLM s on tasks relating to the detection of mental illness symptoms or mental illnesses from social media. We focus on several unique contributions:
  • We test models never used before in this context, including Gemini 1.0 Pro, GPT-4/GPT-4-Turbo/GPT-4o/GPT-4o mini, Mistral NeMo/Medium/Large/Large 2, Claude 3/3.5, Mixtral, Gemma 1/2, Phi-3, Llama 3/3.1, and Qwen 2.
  • We conduct an investigation of published human-annotated mental health datasets, both private and public.
  • We conduct extensive experiments with various prompting methods, including FS prompting and severity prompting.
  • We perform unique experiments such as temperature variation tests, filtering strategies, and chain-of-thought prompting to explore model performance and interpretability.
  • We provide comprehensive details about our methodology, including all prompts, to ensure full transparency and facilitate reproducibility.
  • We provide a detailed list of drawbacks, and challenges that currently exist when exploring the applications of LLM s in psychiatry, along with suggestions for future researchers.
The rest of this paper is organized as follows. Section 2 details our methodology implementation, including the experimental design, datasets, models, prompt templates, evaluation metrics, and parsing of model outputs. Section 3 presents and discusses the results of our experiments, analyzing the performance of various LLM s across the three tasks. Section 4 delves into additional investigations and experiments conducted alongside our main analysis. Finally, Section 5 concludes the paper by summarizing our findings, discussing the challenges encountered, and proposing potential directions for future research.

### v2

@@ Abstract
Large Language Models (LLMs)have shown promise in various domains, including healthcare, with significant potential to transform mental health applications by enabling scalable and accessible solutions. This study aims to provide a comprehensive evaluation of 33 LLMs, ranging from 2 billion to 405+ billion parameters, in performing key mental health tasks using social media data across six datasets. To our knowledge, this represents the largest-scale systematic evaluation of modern LLMsfor mental health applications. Models such as GPT-4, Llama 3, Claude, Gemma, Gemini, and Phi-3 were assessed for their zero-shot (ZS)and few-shot (FS)capabilities across three tasks: binary disorder detection, disorder severity evaluation, and psychiatric knowledge assessment. Key findings revealed that models like GPT-4 and Llama 3 exhibited superior performance in binary disorder detection, achieving accuracies up to 85% on certain datasets, while FSlearning notably enhanced disorder severity evaluations, reducing the Mean Absolute Error (MAE)by 1.3 points for the Phi-3-mini model. Recent models, such as Llama 3.1 405b, demonstrated exceptional psychiatric knowledge assessment accuracy at 91.2%, while prompt engineering played a crucial role in improving performance across tasks. However, the ethical constraints imposed by many LLMproviders limit their ability to respond to sensitive queries, hampering comprehensive performance evaluations. This work highlights both the capabilities and limitations of LLMs in mental health contexts, offering valuable insights for future applications in psychiatry.
@@ 1 Introduction
Artificial Intelligence (AI)is rapidly transforming the landscape of mental healthcare, offering innovative solutions to address the growing global burden of mental illness. According to the World Health Organization, mental disorders accounted for 5.1% of the global disease burden in 2019, with depressive disorders affecting 280 million people worldwide, and an estimated 703,000 people dying by suicide [world2022world]. The economic consequences of mental health conditions are also significant. In the United States, the cumulative cost of mental health inequities is projected to reach a staggering US\$14 trillion between 2024 and 2040. These costs include direct medical expenses, emergency department utilization, productivity losses, and premature deaths [deloitte2024projected].
@@
The integration of AIin psychiatry has enabled a wide array of applications, spanning from the crucial tasks of early detection and diagnosis to predicting treatment outcomes and providing therapeutic interventions. AImodels have been employed to analyze diverse data modalities, including neuroimaging data such as functional magnetic resonance imaging (fMRI)and electroencephalogram (EEG), physiological signals like electrocardiogram (ECG), genetic information, and demographic data [chen2022modern].
@@
The advent of Natural Language Processing (NLP)in mental health research opened up new possibilities. Early NLPapplications focused on analyzing text from clinical interviews and self-report questionnaires. These applications used techniques like the Linguistic Inquiry and Word Count to quantify linguistic patterns associated with various mental health conditions [le2021machine]. The emergence of transformer-based models, such as Bidirectional Encoder Representations from Transformers (BERT)[devlin2018bert], marked a significant turning point in NLPfor mental health. These models, trained on massive amounts of text data, can automatically learn complex linguistic representations, eliminating the need for manual feature engineering. This has led to substantial improvements in tasks like sentiment analysis, suicide risk assessment, and the detection of mental health conditions from social media posts [ji2021mentalbert].
@@
The advancement to LLMs marks a significant leap forward in the field of NLP. LLMs, such as OpenAI’s GPT-4 [achiam2023gpt], Google’s Gemini [team2023gemini, reid2024gemini], Meta’s Llama [touvron2023Llama], and many others have been trained on vast corpora of text data from diverse sources across the internet, encompassing billions of words and a wide range of human knowledge. This extensive training enables LLMs to understand and generate human-like text with a remarkable degree of fluency and coherence.
@@
Given their extensive training, LLMs have immense potential in the realm of mental health and psychiatry. They can be employed for initial consultations, providing patients with preliminary information about mental health disorders and guiding them toward appropriate care pathways. LLMs can quickly analyze transcripts from patient sessions, offering initial insights and supporting medical practitioners in making more informed decisions. Their ability to process large volumes of patient history and previous medication data can streamline administrative tasks, ensuring that practitioners have quick access to comprehensive patient profiles. Additionally, LLMs can serve as virtual counselors, offering support and guidance to individuals experiencing poor mental states, thus extending the reach of mental health services to those who may not have immediate access to professional help. These models can also assist in monitoring social media for signs of mental distress, enabling timely interventions and potentially saving lives. As these technologies continue to evolve, their integration into mental healthcare promises to enhance the efficiency, accessibility, and quality of care provided to patients [obradovich2024opportunities].
Previous research on LLMs in mental health has explored five primary directions (as detailed in Section 2): 1) Initial evaluations of GPT-3.5-Turbo’s capabilities in detecting depression (86% F1) and suicide risk (37% F1) from social media [lamichhane2023evaluation, amin2023will], 2) Fine-tuning smaller models like Llama 2 and FLAN-T5 to match specialized architectures such as Mental-RoBERTa [yang2024mentaLlama, xu2024mental], 3) Developing chatbots for patient simulations [chen2023llm] and systems integrating clinical guidelines like Diagnostic and Statistical Manual of Mental Disorders-V (DSM-V)[qin2023read], 4) Creating evaluation benchmarks like PsyEval for standardized testing [jin2023psyeval], and 5) Literature reviews synthesizing ethical and practical considerations [obradovich2024opportunities, hua2024large]. While these studies demonstrate progress, recent analyses note most focus on older models like GPT-3.5 rather than modern architectures, with limited transparency in prompt engineering methodologies [omar2024applications].
While the potential of LLMs in mental health care is evident, several research gaps remain to be addressed. Firstly, existing evaluations of LLMs in mental health applications are limited in scope, with most studies testing only few models. This narrow focus restricts understanding of how performance varies across architectures, scales, and training approaches, making it difficult to determine which designs are most effective for real-world clinical use. Secondly, there is a lack of focus on prompt engineering to investigate the effect of different prompts on disorder detection tasks. While a few studies have touched upon this aspect, a more systematic and comprehensive exploration of prompt engineering techniques is needed to optimize LLMperformance in this domain. Thirdly, many studies have overlooked smaller yet important factors such as LLMvariation across runs and the effect of small prompt modifications across models. Additionally, there is often a lack of full transparency in reporting methodologies, even when code is provided, hindering reproducibility and further research.
To address these limitations, our study makes the following advances:
  • Comprehensive Model Evaluation: First study to evaluate 33 cutting-edge LLMs (including GPT-4o, Llama 3.1 405B, Claude 3.5, and Phi-3) across parameter scales from 2B to 405B+ - significantly expanding beyond prior work limited to GPT-3.5/4 and Llama 2.
  • Extensive Prompt Engineering: Systematic analysis of ZSvs. FSperformance with severity scoring templates, demonstrating MAEreductions up to 1.3 points through prompt optimization (e.g., in Phi-3-mini).
  • Methodology Transparency: Full disclosure of all prompt templates used across 6 datasets with temperature/seed controls - enabling exact reproducibility missing in previous studies.
  • Dataset Investigation: Comprehensive investigation of published human-annotated mental health datasets, both private and public, ensuring the quality and reliability of data used in experiments by prioritizing datasets labeled by experts or trained volunteers.
  • Practical Challenges Identification: Identified key barriers including high costs (GPT-4/Claude), Application Programming Interface (API)restrictions (Gemini), and prompt sensitivity issues. Suggested future approaches: CoT prompting, RAG integration, and fine-tuning on clinical data.
This paper is organized as follows: Section 2 discusses the related work and usage of LLMs in the mental illness domain. Section 3 outlines the methodology employed in this study, including the experimental design, the datasets and models used, and the creation of prompt templates. Section 4 presents the results of our experiments, analyzing the performance of various LLMs on three key tasks: binary disorder classification, disorder severity evaluation, and psychiatric knowledge assessment. This section begins with an examination of performance variability across models. Section 5 discusses the broader implications of these findings, focusing on model selection considerations, the role of prompt engineering, and the balance between model safety and accuracy in mental health applications. Section 6 details supplementary experiments conducted alongside the main analysis, such as exploring the impact of repeating instructions in prompts and a knowledge reinforcement approach. Section 7.1 acknowledges the limitations of the current study, including dataset constraints and model-related challenges. Section 7.2 suggests potential directions for future research in this area, such as exploring different model architectures and expanding to other mental health tasks. Section 8 concludes the paper.
2 Related Works
Research on LLMs in mental health has rapidly expanded across several directions. Current work in the literature focuses on five main areas: basic evaluation of existing models (Sec. 2.1), performance enhancement through fine-tuning (Sec. 2.2), development of practical applications including data augmentation and chatbots (Sec. 2.3), creation of standardized benchmarks (Sec. 2.4), and systematic literature reviews (Sec. 2.5). While these studies demonstrate promising results, they also highlight the need for extensive validation before clinical deployment.
@@ 2.1 Evaluating LLMs on Mental Health Tasks
One of the earliest works in this domain is the study by [lamichhane2023evaluation], which presents a simple evaluation of GPT-3.5-Turbo’s ZSbinary classification performance on stress, depression, and suicidality detection tasks using social media posts. The resulting F1 scores (73% for stress, 86% for depression, and 37% for suicide), were particularly bad in the suicidality severity prediction task. The authors of [amin2023will] extend this evaluation by assessing GPT-3.5-Turbo’s text classification abilities on a broader range of affective computing problems, including big-five personality prediction and sentiment analysis, in addition to suicide tendency detection. Their findings reveal that while GPT-3.5-Turbo provided decent results, comparable to Word2Vec and bag-of-words baselines, it did not outperform the fine-tuned RoBERTa model on the given tasks. However, the authors did not prompt GPT-3.5-Turbo through an API, they instead did it manually using the web interface, only resetting the context every 25 prompts, which might have affected their results.
@@ 2.2 Fine-tuning LLMs for mental health tasks
Building upon the evaluation studies, [yang2024mentaLlama] and [xu2024mental] explore fine-tuning LLMs for improved performance in mental health tasks. [yang2024mentaLlama] introduce the MentaLlama model, which is fine-tuned on the IMHI dataset, a large dataset of social media posts labeled with mental health conditions and explanations. The authors created the IMHI dataset by taking multiple existing mental health datasets and generating additional explanatory labels using GPT-3.5-Turbo. The MentaLlama-chat-13B model surpasses or approaches state-of-the-art discriminative methods in prediction correctness on 7 out of 10 test sets and generates explanations on par with GPT-3.5-Turbo thus providing interpretability.
The authors of [xu2024mental] conducted a comprehensive evaluation of multiple LLMs, including Alpaca, Alpaca-LoRA, FLAN-T5, Llama 2, GPT-3.5-Turbo, and GPT-4, on various mental health prediction tasks using online text data. Their work explored ZSprompting, FSprompting, and instruction fine-tuning techniques. The authors fine-tuned Alpaca and FLAN-T5 on six mental health prediction tasks across four datasets from Reddit. Results show that instruction fine-tuning significantly improves the performance of LLMs on these tasks. The fine-tuned models, Mental-Alpaca and Mental-FLAN-T5, outperformed much larger models like GPT-3.5-Turbo and GPT-4 on specific tasks, and performed on par with the state-of-the-art task-specific model Mental-RoBERTa [ji2021mentalbert]. The study also includes an exploratory case study on these models’ reasoning capability, which further suggests both the promising future and the important limitations of LLMs.
@@ 2.3 LLMs for Data Augmentation and Chatbot Development
[qin2023read], [bucur2023utilizing], and [chen2023llm] explore the application of LLMs beyond prediction tasks, focusing on data augmentation and chatbot development. [qin2023read] Developed Chat-Diagnose, an explainable and interactive LLM-augmented system for depression detection in social media. Chat-Diagnose incorporates a tweet selector to filter excessive posts, an image descriptor to convert images into text, and professional diagnostic criteria (DSM-V) to guide the diagnostic process. The system utilizes the chain-of-thought (CoT)technique to provide explanations and diagnostic evidence. In the full-training setting, it leverages answer heuristics from a traditional depression detection model. Chat-Diagnose achieves state-of-the-art (SoTA)performance in various settings, including ZS, FS, and full-training scenarios.
[bucur2023utilizing] investigates the use of GPT-3.5-Turbo for generating synthetic Reddit posts simulating depression symptoms from the Beck Depression Inventory (BDI)-II questionnaire, aiming to enhance semantic search capabilities. The study finds that using original BDI-II responses as queries is more effective, suggesting that the generated data might be too specific for this task.
[chen2023llm] focuses on developing chatbots that simulate psychiatrists and patients in clinical diagnosis scenarios, specifically for depressive disorders. The study follows a three-phase design, involving collaboration with psychiatrists, experimental studies, and evaluations with real psychiatrists and patients. The development process emphasizes an iterative refinement of prompt design and evaluation metrics based on feedback from both psychiatrists and patients. A key finding of the study is the comparison between real and simulated psychiatrists, revealing differences in questioning strategies and empathy behaviors.
@@ 2.4 Benchmarks for Evaluating LLMs in Psychiatry
In contrast to the aforementioned studies, the authors of [jin2023psyeval] took a different approach. Instead of evaluating existing LLMs on established datasets, they sought to create a comprehensive benchmark specifically designed to assess the capabilities of LLMs in the mental health domain. PsyEval (the benchmark) includes tasks such as mental health question-answering, where LLMs are evaluated on their ability to provide accurate and informative responses to queries related to mental health conditions, symptoms, and treatments. The benchmark also features tasks focused on diagnosis prediction, both using data from social media or simulated dialogues between patients and clinicians. Additionally, PsyEval incorporates tasks that assess LLMs’ ability to provide empathetic and safe psychological counseling in simulated conversations.
@@ 2.5 Literature Reviews
In addition to the contribution papers mentioned above, several surveys have been conducted to summarize and analyze the landscape of LLMs in mental health care. The authors of [obradovich2024opportunities] provide a comprehensive overview of the opportunities and risks associated with LLMs in psychiatry, discussing their potential to enhance mental health care through improved diagnostic accuracy, personalized care, and streamlined administrative processes. The authors highlight the ability of LLMs to efficiently analyze patient data, summarize therapy sessions, and aid in complex diagnostic problem-solving. They also discuss the potential of LLMs to automate certain managerial decisions within hospital systems, such as personnel schedules and equipment needs. However, the authors also acknowledge the risks associated with LLMs, such as labor substitution, the potential for reduced human-to-human socialization, and the amplification of existing biases. They conclude by advocating for the development of pragmatic frameworks, including red-teaming and multi-stakeholder-oriented safety and ethical guidelines, to ensure the safe and responsible deployment of LLMs in psychiatry.
@@
The authors of [omar2024applications] conducted a systematic review of the applications of LLMs, such as GPT-3.5-Turbo, in the field of psychiatry. The review identified 16 studies that directly examined the use of LLMs in psychiatry, with applications ranging from clinical reasoning and social media analysis to education. The authors found that LLMs like GPT-3.5-Turbo and GPT-4 showed promise in assisting with tasks such as diagnosing mental health issues, managing depression, evaluating suicide risk, and supporting education in the field. However, the review also highlighted limitations, such as difficulties with complex cases and potential underestimation of suicide risks.
@@
The article by [hua2024large] presents a scoping review of LLMs in mental health care, focusing on their applications and outcomes. The review identified 34 publications that met their inclusion criteria, including studies compiling new datasets, studies focused on model development, studies focused on employing and evaluating LLMs, and studies examining the potential opportunities, risks and ethical considerations associated with utilizing LLMs in a psychiatric context. The authors provide a detailed overview of the models and datasets used in these studies. They found that most studies utilized GPT-3.5 or GPT-4, with some fine-tuning on mental health-specific datasets like MentaLlama and Mental-LLM. The datasets used in these studies were primarily sourced from social media platforms and online mental health forums, with a few utilizing clinical data or simulated conversations. The authors also discuss the key challenges associated with LLMs, such as data availability and reliability, effective handling of mental states, and appropriate evaluation methods.